# BILBO: BILevel Bayesian Optimization

### v1

@@ Abstract
Machine Learning, Bayesian Optimization, Bilevel
@@ 1 Introduction
Figure 1: Example of bilevel optimization with upper-level variable $\mathbf{x}$ and lower-level variable $\mathbf{z}$. The bilevel solution (blue cross) is constrained by lower-level solutions (yellow line) and differs from the non-bilevel solution (red cross).
@@
  • Function query strategy that includes the uncertainty of estimated lower-level solution and a conditional reassignment of the query point for more effective lower-level objective evaluation. This balances exploiting estimated lower-level solutions with exploring to refine these estimates. We show that our strategy leads to an instantaneous regret bound on the query.
We further show that BILBO has sublinear cumulative and simple regret bounds for commonly used kernels, and demonstrate its effectiveness empirically on several synthetic and real-world problems. The code will be released upon publication. A notation table is given in Appendix A.
@@ 2 Related Work
Constrained Bayesian optimization. Several constrained BO algorithms have been proposed (Gelbart et al., 2014; Hernández-Lobato et al., 2016). In particular, Xu et al. (2023) and Nguyen et al. (2023) both introduced confidence-bound based optimistic estimations of the feasible set, with the former providing an infeasibility declaration scheme and the latter including a function query strategy for decoupled settings. These feasible set estimations guide sampling toward probable feasible points, improving sample efficiency. Compared to constrained optimization, bilevel optimization presents additional challenges due to the need to optimize a separate lower-level problem, where the optimal solutions are unknown and often estimated suboptimally.
Comparison to other optimization problems. While some optimization problems, such as robust optimization (Bogunovic et al., 2018) and composite objectives optimization (Li & Scarlett, 2024), respectively involve an additional random variable and composite objective function, they remain single-level problems. In contrast, bilevel optimization involves a two-level hierarchical structure, where the upper-level is constrained by lower-level solutions, making it fundamentally different from these other settings.
@@ 3 Preliminaries
Bayesian optimization. Given a prior distribution $P(h)$ and likelihood function $P(\mathcal{D}_{h_{t},t}|h)$, the posterior distribution $P(h|\mathcal{D}_{h_{t},t})$ can be calculated via Bayes’ theorem. The prior distribution is often represented by a GP and the likelihood function is defined by the choice of GP kernel and hyperparameters. The posterior distribution is also a surrogate model for $h$. The point which maximizes an acquisition function $a_{h}(\mathbf{xz})$ is selected as the next point to evaluate function $h$ at. A popular acquisition strategy is based on confidence bounds (Srinivas et al., 2010).
@@
$r_{F}(\mathbf{x}_{t},\mathbf{z}_{t})$ $\triangleq F(\mathbf{x}^{*},\mathbf{z}^{*})-F(\mathbf{x}_{t},\mathbf{z}_{t}),$ (3.5)
@@ 4 Bilevel Bayesian optimization
For some small $\delta>0$, with probability at least $1-\delta$,
@@
$\forall(\mathbf{x},\mathbf{z})\in\mathcal{S}^{+}_{t},c\in\mathcal{C}_{\text{up}}\cup\mathcal{C}_{\text{lo}}$, the constraint regrets are upper bounded,
@@ 4.1 Trusted set of optimal lower-level solutions
$\forall(\mathbf{x},\mathbf{z})\in\mathcal{P}^{+}_{t}$, the lower-level objective regret is upper bounded,
@@ 4.3 Function query
Following the function query selection in Definition 4.7 and reassignment of query point in Section 4.3, the instantaneous regret for the query point $(\mathbf{x}_{t},\mathbf{z}_{t})$ at time $t\geq 1$ is upper bounded by,
@@ 4.4 Regret bound
With probability at least $1-\delta$, $T\geq 1$, and $\beta_{T}\triangleq 2\log(|\mathcal{F}||\mathcal{X}||\mathcal{Z}|T^{2}\pi^{2}/6\delta)$, the estimator $(\hat{\mathbf{x}}_{T},\hat{\mathbf{z}}_{T})$, defined in Equation 4.9, has a simple regret bound of
@@ 5 Experiments
Figure 3: BraninHoo+GoldsteinPrice experiment details. (a) Upper-level objective, Branin-Hoo. (b) Lower-level objective, Goldstein-Price. (c) BILBO’s upper-level estimate. (d) BILBO’s lower-level estimate. (e-f) BILBO’s iterative queries.
@@ 5.1 Synthetic problems
BraninHoo+GoldsteinPrice has the Branin-Hoo function as upper-level objective $F$ and the Goldstein-Price function as the lower level objective $f$ (Picheny et al., 2013). Both functions are non-convex and multimodal. The dimensions $d_{\mathcal{X}}$ and $d_{\mathcal{Z}}$ are both 1, which facilitates visualization of the models and queries. Both dimensions were discretized into 100 points. The Branin-Hoo function has 3 optimal points, but there is only 1 optimal bilevel solution when constrained by lower-level Goldstein-Price optimal solutions.
@@ 6 Future Work
Adaptive discretization (Shekhar & Javidi, 2017) may also reduce computational complexity by reducing the effective dimension of the explored space. Discretization strategies could be integrated with trusted sets, for example concentrating the discretizations within the trusted sets. Approximate surrogate models (Calandriello et al., 2019) is another possible direction for scalability while preserving confidence bound estimates. The theoretical work presented in this paper could be extended to approximate surrogate models.
@@ Appendix C Proofs
$r_{F}(\mathbf{x}_{t},\mathbf{z}_{t})$ $\triangleq F(\mathbf{x}^{*},\mathbf{z}^{*})-F(\mathbf{x}_{t},\mathbf{z}_{t})$ from Equation 3.5
$\leq u_{F,t}(\mathbf{x}^{*},\mathbf{z}^{*})-l_{F,t}(\mathbf{x}_{t},\mathbf{z}_{t})$ from Corollary 4.2
@@ D.4 Energy market
Let $\mathbf{x}\triangleq[\mathbf{x}_{1},\mathbf{x}_{2}]$, where $\mathbf{x}_{1}$ denotes a price to bid and $\mathbf{x}_{2}$ denotes a quantity in MW to supply at bid price. $\mathbf{x}_{1}\in(0.01,0.5),\mathbf{x}_{2}\in(200,500)$. We simulate a network with 3 generators that has to fulfill an estimated demand schedule for 3 periods. The generators’ parameters are given in Table 2, where $\mathbf{z}\triangleq[\mathbf{z}_{1},\mathbf{z}_{2}]$ are the lower-level variables. $\mathbf{z}_{1}\in(0.0,0.2),\mathbf{z}_{2}\in(0.5,1.5)$. These two variables were selected as a proxy for auxiliary concerns such as efficiency and maintainence costs, on top of operational costs.
@@
Table 2: Parameters input into PyPSA generator. ‘max p factor‘ refer to ‘p_max_pu‘, the maximum power at a snapshot given as a fraction of nominal power.
@@ D.6 Computational resources
The experiments in this paper were done on a computer with AMD Ryzen 7 5700X 8-Core Processor and 64 GB of RAM.

### v2

@@ Abstract
machine learning,Bayesian optimization,bilevel optimization,bilevel Bayesian optimization
@@ 1 Introduction
Figure 1: Example of bilevel optimization with upper-level variable $\mathbf{x}$ and lower-level variable $\mathbf{z}$. The bilevel solution (blue cross) is constrained by lower-level (LL) solutions (yellow line) and differs from the non-bilevel solution (red cross).
@@
  • Function query strategy that includes the uncertainty of the estimated lower-level solution and a conditional reassignment of the query point for more effective lower-level objective evaluation. This balances exploiting estimated lower-level solutions with exploring to refine these estimates. We show that our strategy leads to an instantaneous regret bound on the query.
We further show that BILBO has sublinear cumulative and simple regret bounds for commonly used kernels, and demonstrate its effectiveness empirically on several synthetic and real-world problems. The code is provided in https://github.com/chewwt/bilbo/, and a notation table is in Appendix A.
@@ 2 Related Work
Constrained Bayesian optimization. Several constrained BO algorithms have been proposed (Gardner et al., 2014; Gelbart et al., 2014; Hernández-Lobato et al., 2016; Eriksson & Poloczek, 2021; Takeno et al., 2022). In particular, Xu et al. (2023) and Nguyen et al. (2023) both introduced confidence-bound based optimistic estimations of the feasible set, with the former providing an infeasibility declaration scheme and the latter including a function query strategy for decoupled settings. These feasible set estimations guide sampling toward probable feasible points, improving sample efficiency. Compared to constrained optimization, bilevel optimization presents additional challenges due to the need to optimize a separate lower-level problem, where the optimal solutions are unknown and often estimated suboptimally.
Comparison to other optimization problems. While some optimization problems, such as robust optimization (Bogunovic et al., 2018; Kirschner et al., 2020) and composite objectives optimization (Astudillo & Frazier, 2019; Li & Scarlett, 2024), respectively involve an additional random variable and composite objective function, they remain single-level problems. In contrast, bilevel optimization involves a two-level hierarchical structure, where the upper-level is constrained by lower-level solutions, making it fundamentally different from these other settings.
@@ 3 Preliminaries
Bayesian optimization. Given a prior distribution $P(h)$ and likelihood function $P(\mathcal{D}_{h_{t},t}|h)$, the posterior distribution $P(h|\mathcal{D}_{h_{t},t})$ can be calculated via Bayes’ theorem. The prior distribution is often represented by a GP and the likelihood function is defined by the choice of GP kernel and hyperparameters. The posterior distribution is also a surrogate model for $h$. The point which maximizes an acquisition function $a_{h}(\mathbf{xz})$ is selected as the next point to evaluate function $h$ at (Brochu et al., 2010; Frazier, 2018; Garnett, 2023). Popular acquisition strategies include information-theoretic ones (Hennig & Schuler, 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017) and upper confidence bounds (Srinivas et al., 2010).
@@
$r_{F}(\mathbf{x}_{t},\mathbf{z}_{t})$ $\triangleq\max(0,F(\mathbf{x}^{*},\mathbf{z}^{*})-F(\mathbf{x}_{t},\mathbf{z}_{t})),$ (3.5)
@@ 4 Bilevel Bayesian optimization
For some small $\delta>0$, with probability at least $1-\delta$, $\forall\mathbf{x}\in\mathcal{X},\mathbf{z}\in\mathcal{Z},h\in\mathcal{F}$, and $t\geq 1$,
@@
Let $\delta\in(0,1)$, with probability at least $1-\delta$, $\forall(\mathbf{x},\mathbf{z})\in\mathcal{S}^{+}_{t},c\in\mathcal{C}_{\text{up}}\cup\mathcal{C}_{\text{lo}}$, the constraint regrets are upper bounded,
@@ 4.1 Trusted set of optimal lower-level solutions
Let $\delta\in(0,1)$, with probability at least $1-\delta$, $\forall(\mathbf{x},\mathbf{z})\in\mathcal{P}^{+}_{t}$, the lower-level objective regret is upper bounded,
@@ 4.3 Function query
Let $\delta\in(0,1)$, with probability at least $1-\delta$, following the function query selection in Definition 4.7 and reassignment of query point in Section 4.3, the instantaneous regret for the query point $(\mathbf{x}_{t},\mathbf{z}_{t})$ at time $t\geq 1$ is upper bounded by,
@@ 4.4 Regret bound
Let $\delta\in(0,1)$, with probability at least $1-\delta$, $T\geq 1$, and $\beta_{T}\triangleq 2\log(|\mathcal{F}||\mathcal{X}||\mathcal{Z}|T^{2}\pi^{2}/6\delta)$, the estimator $(\hat{\mathbf{x}}_{T},\hat{\mathbf{z}}_{T})$, defined in Equation 4.9, has a simple regret bound of
@@ 5 Experiments
Figure 3: BraninHoo+GoldsteinPrice experiment details. LL refers to lower-level. (a) Upper-level objective, Branin-Hoo. (b) Lower-level objective, Goldstein-Price. (c) BILBO’s upper-level estimate. (d) BILBO’s lower-level estimate. (e-f) BILBO’s iterative queries.
@@ 5.1 Synthetic problems
BraninHoo+GoldsteinPrice has the Branin-Hoo function as upper-level objective $F$ and the Goldstein-Price function as the lower-level objective $f$ (Picheny et al., 2013). Both functions are non-convex and multimodal. The dimensions $d_{\mathcal{X}}$ and $d_{\mathcal{Z}}$ are both 1, which facilitates visualization of the models and queries. Both dimensions were discretized into 100 points. The Branin-Hoo function has 3 optimal points, but there is only 1 optimal bilevel solution when constrained by lower-level Goldstein-Price optimal solutions.
@@ 5.2 Real-world problems
Additional discussions on the computational complexity and time efficiency of BILBO can be found in Appendix E.
@@ 6 Future Work
Adaptive discretization (Shekhar & Javidi, 2018) may also reduce computational complexity by reducing the effective dimension of the explored space. Discretization strategies could be integrated with trusted sets, for example concentrating the discretizations within the trusted sets. Approximate surrogate models (Calandriello et al., 2019) is another possible direction for scalability while preserving confidence bound estimates. The theoretical work presented in this paper could be extended to approximate surrogate models.
@@
Future work can extend our discretized implementation to continuous domains, which should be feasible with a scalable trusted set representation. Our theoretical results can also generalize to continuous settings with minor modifications to $\beta_{t}$ and additional assumptions, following Chowdhury & Gopalan (2017).
@@ 7 Conclusion
Acknowledgments
This research is supported by the National Research Foundation (NRF), Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3333S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre.
@@ Appendix C Proofs
$r_{F}(\mathbf{x}_{t},\mathbf{z}_{t})$ $\triangleq\max(0,F(\mathbf{x}^{*},\mathbf{z}^{*})-F(\mathbf{x}_{t},\mathbf{z}_{t}))$ from Equation 3.5
$\leq\max(0,u_{F,t}(\mathbf{x}^{*},\mathbf{z}^{*})-l_{F,t}(\mathbf{x}_{t},\mathbf{z}_{t}))$ from Corollary 4.2
@@ D.4 Energy market
Let $\mathbf{x}\triangleq[\mathbf{x}_{1},\mathbf{x}_{2}]$, where $\mathbf{x}_{1}$ denotes a price to bid and $\mathbf{x}_{2}$ denotes a quantity in MW to supply at bid price. $\mathbf{x}_{1}\in(0.01,0.5),\mathbf{x}_{2}\in(200,500)$. We simulate a network with 3 generators that has to fulfill an estimated demand schedule for 3 periods. The generators’ parameters are given in Table 2, where $\mathbf{z}\triangleq[\mathbf{z}_{1},\mathbf{z}_{2}]$ are the lower-level variables. $\mathbf{z}_{1}\in(0.0,0.2),\mathbf{z}_{2}\in(0.5,1.5)$. These two variables were selected as a proxy for auxiliary concerns such as efficiency and maintenance costs, on top of operational costs.
@@
Table 2: Parameters input into PyPSA generator. ‘max p factor’ refer to ‘p_max_pu’, the maximum power at a snapshot given as a fraction of nominal power.
@@ D.6 Computational resources
The experiments in this paper were done on a computer with AMD Ryzen 7 5700X 8-Core Processor and 64 GB of RAM, unless otherwise specified.
Appendix E Complexity and efficiency of BILBO
E.1 Computational complexity of BILBO
For a discretized implementation, the computational complexity of BILBO is affected by the computational complexity of:
  • Gaussian processes, $\mathcal{O}(n^{3})$,
  • Updating trusted sets, $\mathcal{O}(|\mathcal{F}|c)$,
  • Selecting function query, $\mathcal{O}(|\mathcal{F}|c)$,
  • Optimizing the acquisition function, $\mathcal{O}(c)$,
where $n$ is the number of observations, $c$ is the number of discretized points, and $|\mathcal{F}|$ is the number of blackbox functions.
In a uniform grid discretization, which is used in our implementation, if each dimension is divided into $m$ points, then the cardinality is $c=m^{d}$, where $d$ is the number of dimensions. Thus, dimensionality $d$ exponentially affects computational complexity when using uniform grid discretization. Adaptive discretization may be able to mitigate the exponential factor of dimensionality, where effective dimension $d_{\text{eff}}\ll d$.
E.2 Time efficiency of BILBO
While our experiments have shown that BILBO is more sample efficient than nested methods, BILBO does require more computational cost to update trusted sets and select function queries and query points. In the presence of inexpensive lower-level evaluations, BILBO’s time efficiency can be lower than that of a nested method. However, do note that BILBO’s motivation is for settings with expensive blackbox evaluations, where evaluations can be real-world experiments or simulators that are costly or slow.
In addition, BILBO has advantages in scenarios with noisy observations or multiple lower-level solutions. Nested methods only solve for one solution in each lower-level optimization, and it can be suboptimal in these scenarios. On the other hand, BILBO manages the uncertainty of lower-level estimates in a principled way and allows for multiple lower-level estimates, possibly providing better lower-level estimates to reduce regret more effectively even if each iteration takes more time.
In terms of wall-clock time, on a Mac Studio with M2 Ultra over 5 runs and 40 seconds total runtime, for the 2-dimensional BraninHoo+GoldsteinPrice experiment, the average time per BILBO iteration is 0.131s, which is about 1.5 times slower than TrustedRand and about 26 times slower than Nested. Figure 7 shows how regret decreases over wall-clock time. We observed that while the regret for Nested is smaller than BILBO in the initial 5 seconds, Nested’s regret quickly plateaus due to suboptimal lower-level estimates of the multimodal lower-level objective. BILBO outperforms Nested subsequently as BILBO converges to a more optimal solution.
Figure 7: Regret against wall-clock time for the BraninHoo+GoldsteinPrice experiment